# Improved particle-flow event reconstruction with scalable neural networks for current and future particle detectors

### v5

@@ 1 Introduction
There has been considerable interest in and development of machine learning (ML)-based reconstruction methods, including for PF reconstruction. In order to train a model that reconstructs events consisting of a variable number of particles, a recipe for a loss function based on attractive and repulsive potentials was given in Ref. [17]. In Ref. [18], a version of this loss was used together with a graph neural network (GNN)-based approach to reconstruct events with high particle multiplicity accurately. This approach was successfully applied in the CMS experiment [19, 20]. Recently, Ref. [21], demonstrated that a network architecture based on learning a hypergraph structure can improve jet reconstruction. In parallel, clustering using ML has been demonstrated for high-granularity calorimeter reconstruction [22].
@@ 2.1 Loss function
The input to the model is a set of detector elements: charged tracks and calorimeter clusters (or alternatively, raw calorimeter hits), each described by a feature vector $\mathbf{x}\in X$. The input features for the tracks consist of the track $p_{\mathrm{T}}$, the pseudorapidity of the particle momentum (track tangent) $\eta$, the azimuthal angle of the particle momentum $\phi$, the track goodness-of-fit $\chi^{2}$ and the number of degrees of freedom $N_{\mathrm{dof}}$, the slope of the track $\tan{\lambda}=p_{z}/\sqrt{p_{x}^{2}+p_{y}^{2}}$, the signed impact parameter $D_{0}$ with respect to the origin $(0,0,0)$ at interaction point in the $xy$ plane, the curvature of the track $\Omega=1/R$ defined via the track radius $R$, the position of the track along the $z$-axis $Z_{0}$ with respect to the origin [29]. For calorimeter clusters, the features consist of transverse energy $E_{\mathrm{T}}$, $\eta$, $\phi$, electromagnetic calorimeter energy $E_{\mathrm{ECAL}}$, hadronic calorimeter energy $E_{\mathrm{HCAL}}$, Cartesian spatial coordinates $x$, $y$, $z$; the number of hits and the cluster size, measured in standard deviations of the hit sets in $x$, $y$, and $z$.
The target of the model is a set of stable particles such as charged and neutral pions, photons, electrons and muons, each described by a particle type, charge, and continuous four-momentum values, combined to form a target feature vector $\mathbf{y}\in Y$. The goal of the model is, for each event, to reconstruct the set $Y$, given the set $X$, i.e. $\Phi(X)\rightarrow Y^{\prime}\simeq Y$. To make the set-to-set translation tractable with ML, and to be able to treat particle multiplicities in the range of $|X|,|Y|\simeq 10^{4}$ with a single model pass, each target particle is associated with a single unique input element.
@@ 3 Results and discussion
An extensive HPO was performed for both the GNN- and kernel-based transformer model using the JURECA supercomputer [52]. JURECA is a pre-exascale modular supercomputer operated by Jülich Supercomputing Centre at Forschungszentrum Jülich. The system consists of a flexible Data Centric (DC) module, based on the Atos BullSequana XH2000. It has, among others, 192 accelerated compute nodes with four NVIDIA A100 GPUs and two AMD EPYC 7742 CPUs each. We employ a similar approach to HPO as in Ref. [53], using Bayesian Optimization in combination with the ASHA [54] algorithm. The optimization was performed in a distributed manner on 96 GPUs spread over 24 compute nodes, consuming roughly 7000 and 5000 GPU hours for the optimization of the kernel-based transformer and the GNN model respectively. In Tables 2 and 2, we show the search spaces used for the HPO of the GNN and kernel-based transformer models respectively, while Tables 4 and 4 show the best HPs found for each model. Both search spaces are relatively large, each containing 10 different HPs with between three and six possible values. Tables 5 and 6 contain descriptions of the HPs in the two search spaces.
Table 1: Search space used for HPO of the GNN model
Table 2: Search space used for HPO of the transformer model
Hyperparameter Search space
lr {$10^{-4}$, $10^{-3}$, $10^{-2}$}
batch_multiplier {10, 20, 40}
num_graph_layers_id {1, 2, 3, 4, 5, 6}
num_graph_layers_reg {1, 2, 3, 4, 5, 6}
bin_size {16, 32, 64, 128, 256}
output_dim {64, 128, 256}
ffn_dist_hidden_dim {64, 128, 256}
ffn_dist_num_layers {1, 2, 3, 4, 5}
out_hidden_dim {64, 128, 256, 512}
out_num_layers {1, 2, 3, 4, 5}
Hyperparameter Search space
lr {$10^{-5}$, $10^{-4}$, $10^{-3}$}
batch_multiplier {10, 20, 40}
num_layers_encoder {1, 2, 3, 4}
num_layers_decoder_reg {1, 2, 3, 4}
num_layers_decoder_cls {1, 2, 3, 4}
hidden_dim {32, 64, 128}
num_heads {8, 16, 32, 64}
num_random_features {16, 32, 64, 128}
out_hidden_dim {128, 256, 512}
out_num_layers {1, 2, 3, 4}
Table 2: Search space used for HPO of the transformer model
Table 3: Best HPs found for the GNN model
Table 4: Best HPs found for the transformer model
Hyperparameter Search space
lr $10^{-3}$
batch_multiplier 10
num_graph_layers_id 6
num_graph_layers_reg 6
bin_size 256
output_dim 256
ffn_dist_hidden_dim 64
ffn_dist_num_layers 3
out_hidden_dim 128
out_num_layers 1
Hyperparameter Search space
lr $10^{-3}$
batch_multiplier 10
num_layers_encoder 3
num_layers_decoder_reg 2
num_layers_decoder_cls 2
hidden_dim 32
num_heads 8
num_random_features 32
out_hidden_dim 128
out_num_layers 1
Table 4: Best HPs found for the transformer model
Table 5: Descriptions of HPs in the GNN model
Hyperparameter Search space
lr learning rate
batch_multiplier batch size multiplier
num_graph_layers_id number of graph layers for classification
num_graph_layers_reg number of grapg layers for regression
bin_size size of LSH bins
output_dim output dimension of graph convolution layers
ffn_dist_hidden_dim hidden dimension of feed-forward layer
ffn_dist_num_layers number of feed-forward layers
out_hidden_dim hidden dimension of output decoding MLP
out_num_layers number of layers of output decoding MLP
Table 6: Descriptions of HPs in the transformer model
Hyperparameter Search space
lr learning rate
batch_multiplier batch size multiplier
num_layers_encoder number of layers in the encoder
num_layers_decoder_reg number of layers in the regression decoder
num_layers_decoder_cls number of layers in the classification decoder
hidden_dim hidden dimension
num_heads number of trasnformer heads
num_random_features number of random features
out_hidden_dim hidden dimension of output decoding MLP
out_num_layers number of layers of output decoding MLP
@@
Charged hadron efficiency in (a), neutral hadron efficiency in (b), photon efficiency in (c). The efficiencies are parameterized as a function of track $p_{\mathrm{T}}$ or cluster energy. The charged hadron fake rate is shown in (d), the neutral hadron fake rate in (e), the photon fake rate in (f). For efficiency and fake rate, we show the binomial statistical uncertainties from limited samples.
@@
An important factor in the development of scalable ML-based full event reconstruction models is improving the computational throughput in future deployment scenarios. Possible approaches to improve inference throughput could include more efficient model formulations and implementations [56], sparsity and quantization. Therefore, we compare the inference scalability of the rule-based PF implementation on CPU and the proposed scalable GNN implementation on GPU with respect to increasing input multiplicity, with the results summarized in Fig. 10. Inference scaling tests of the GNN model are tested on a small, 8 GB consumer GPU (Nvidia RTX2070S) to emulate a resource-constrained scenario in edge deployment.
@@ Acknowledgements
We are grateful to the reviewers for their careful reading of the manuscript and their comments which significantly improved the clarity of this paper.

### v6

@@ 1 Introduction
There has been considerable interest in and development of machine learning (ML)-based reconstruction methods, including for PF reconstruction. Models for PF reconstruction based on computer vision were investigated in [17]. In order to train a model that reconstructs events consisting of a variable number of particles, a recipe for a loss function based on attractive and repulsive potentials was given in Ref. [18]. In Ref. [19], a version of this loss was used together with a graph neural network (GNN)-based approach to reconstruct events with high particle multiplicity accurately. This approach was successfully applied in the CMS experiment [20, 21]. Recently, Ref. [22], demonstrated that a network architecture based on learning a hypergraph structure can improve jet reconstruction. In parallel, clustering using ML has been demonstrated for high-granularity calorimeter reconstruction [23].
@@ 2.1 Loss function
The input to the model is a set of detector elements: charged tracks and calorimeter clusters (or alternatively, raw calorimeter hits), each described by a feature vector $\mathbf{x}\in X$. The input features for the tracks consist of the track $p_{\mathrm{T}}$, the pseudorapidity of the particle momentum (track tangent) $\eta$, the azimuthal angle of the particle momentum $\phi$, the track goodness-of-fit $\chi^{2}$ and the number of degrees of freedom $N_{\mathrm{dof}}$, the slope of the track $\tan{\lambda}=p_{z}/\sqrt{p_{x}^{2}+p_{y}^{2}}$, the signed impact parameter $D_{0}$ with respect to the origin $(0,0,0)$ at interaction point in the $xy$ plane, the signed curvature of the track $\Omega=\mathrm{sign}(q)/R$ defined via the track radius $R$ and charge $q$, the position of the track along the $z$-axis $Z_{0}$ with respect to the origin [30]. For calorimeter clusters, the features consist of transverse energy $E_{\mathrm{T}}$, $\eta$, $\phi$, electromagnetic calorimeter energy $E_{\mathrm{ECAL}}$, hadronic calorimeter energy $E_{\mathrm{HCAL}}$, Cartesian spatial coordinates $x$, $y$, $z$; the number of hits and the cluster size, measured in standard deviations of the hit sets in $x$, $y$, and $z$.
The target of the model is a set of stable particles such as charged and neutral pions, photons, electrons, and muons, defined through sufficient energy matched to detector hits and with a Pythia8 8 status code 1. This specific choice of target particles represents a reasonable optimization target, but may not be optimal for all cases. The definition of an optimal ground truth for particle reconstruction is left to future work. Each particle is described by a particle type, charge, and continuous four-momentum values, combined to form a target feature vector $\mathbf{y}\in Y$. The goal of the model is, for each event, to reconstruct the set $Y$, given the set $X$, i.e. $\Phi(X)\rightarrow Y^{\prime}\simeq Y$. To make the set-to-set translation tractable with ML, and to be able to treat particle multiplicities in the range of $|X|,|Y|\simeq 10^{4}$ with a single model pass, each target particle is associated with a single unique input element.
@@ 3 Results and discussion
An extensive HPO was performed for both the GNN- and kernel-based transformer model using the JURECA supercomputer [53]. JURECA is a pre-exascale modular supercomputer operated by Jülich Supercomputing Centre at Forschungszentrum Jülich. The system consists of a flexible Data Centric (DC) module, based on the Atos BullSequana XH2000. It has, among others, 192 accelerated compute nodes with four NVIDIA A100 GPUs and two AMD EPYC 7742 CPUs each. We employ a similar approach to HPO as in Ref. [54], using Bayesian Optimization in combination with the ASHA [55] algorithm. The optimization was performed in a distributed manner on 96 GPUs spread over 24 compute nodes, consuming roughly 7000 and 5000 GPU hours for the optimization of the kernel-based transformer and the GNN model respectively. The hyperparameter optimization details can be found in Supplementary Note 1: Hyperparameter optimization.
@@
Charged hadron efficiency in (a), neutral hadron efficiency in (b), photon efficiency in (c). The efficiencies are parameterized as a function of track $p_{\mathrm{T}}$ or cluster energy. The charged hadron fake rate is shown in (d), the neutral hadron fake rate in (e), the photon fake rate in (f). For efficiency and fake rate, we show the binomial statistical uncertainties from limited samples. We show the $p_{\mathrm{T}}$ response of charged hadrons in (g), and the energy response of neutral hadrons in (h) and photons in (i).
@@
An important factor in the development of scalable ML-based full event reconstruction models is improving the computational throughput in future deployment scenarios. Possible approaches to improve inference throughput could include more efficient model formulations and implementations [57], sparsity and quantization. Therefore, we compare the inference scalability of the rule-based PF implementation on CPU and the proposed scalable GNN implementation on GPU with respect to increasing input multiplicity, with the results summarized in Fig. 10. Inference scaling tests of the GNN model are tested on a small, 8 GB consumer GPU (Nvidia RTX2060S) to emulate a resource-constrained scenario in edge deployment.
@@ Acknowledgements
We are grateful to the reviewers and the editor for their careful reading of the manuscript and their comments which significantly improved the clarity of this paper.
We would also like to thank Nilotpal Kakati and Etienne Dreyer for cross-checks of our published code and datasets.